# Resource-Theoretic Hierarchy of Contextuality for General Probabilistic Theories

### v1

@@ 5.2 Relation with previous works on contextuality and GPTs
Other influential sources of inspiration for our work are [47, 51, 52]. As we already mentioned in the introduction, [47] introduces the notion of simplex embeddability as the geometrical criterion to assess whether a GPT system is noncontextual. [^7] Another work that is closely related to [47] is that of Shahandeh [50]. However, the criterion of classicality introduced there additionally requires the simplex to have the same dimension of the GPT system. We do not adopt this approach. For example, it deems the rebit stabilizer theory to be nonclassical. However, the rebit stabilizer theory, in the prepare-and-measure scenario, admits a noncontextual ontological model given by the Spekkens toy theory [66], as has been shown in several works [77, 78, 79]. The criterion is extended to accessible GPT fragments in [51], where the latter correspond to more general mathematical objects than GPTs and characterize generic prepare-and-measure experimental setups. The work of [52] provides an algorithm for testing contextuality in any prepare-and-measure scenario. In particular, if it exists, it returns an explicit noncontextual model for the scenario and, if not, it provides the minimum amount of (depolarizing) noise which would be required until a noncontextual model would become possible. They call this measure of contextuality the robustness of nonclassicality. The latter is related to our notion of error of univalent simulation, but it is different insofar as it requires a specific noise channel (e.g. depolarizing noise channel), whereas the error of univalent simulation does not require that. Moreover, the robustness of nonclassicality is not defined as a measure within a well defined resource theory.
@@
It has been argued in [51] that in realistic experiments GPT systems are generally not appropriate to study contextuality. In particular, the following discrepancy arises: A GPT system obtained from an operational description of the experiment, as in Remark 9, provides states and effects observable in practice in this experiment, but does not limit the in principle implementable ones. However, the preparation equivalences preserved by a preparation-noncontextual ontological model (see Section 2 and Definition 10) are in principle operational equivalences.
When studying preparation contextuality only, this discrepancy can be dealt with. Namely, one can first set-up the effect space $E_{\mathsf{A}}$ in a non-operational way, as a hypothesized set of in principle possible effects. Importantly, one has to identify each actual measurement outcome arising in the experiment with a (distribution over) the elements of $E_{\mathsf{A}}$. Their means specify a subset $\tilde{E}_{\mathsf{A}}\subseteq E_{\mathsf{A}}$ of the so-called accessible effects, which however play no role in the study of preparation contextuality. The state space $\Omega_{\mathsf{A}}$ is the set of equivalence classes of preparation procedures as in Remark 9, but only with respect to in principle operational equivalence. That is, states that are indistinguishable by accessible effects, but could be distinguished by a hypothesized effect in $E_{\mathsf{A}}\setminus\tilde{E}_{\mathsf{A}}$ ought not be identified in $\Omega_{\mathsf{A}}$. With this set-up, one can study preparation contextuality as we discuss shortly in Section 3.3 in terms of the GPT system $(\Omega_{\mathsf{A}},E_{\mathsf{A}})$.
@@
$\mathcal{P}=\{0,1\},\quad$ =$\mathcal{M}=\{0\},\quad$ $\mathcal{D}=\{0,1\},$ $\mathcal{E}_{P}=\emptyset,$ $\mathcal{E}_{M}=\emptyset,$ (75)
$\mathcal{P}=\{0,1,2\},\quad$ =$\mathcal{M}=\{0\},\quad$ $\mathcal{D}=\{0,1,2\},\quad$ $\mathcal{E}_{P}=\emptyset,\quad$ $\mathcal{E}_{M}=\emptyset.$ (76)
@@
$p(\tilde{k}|\tilde{j}=1,\tilde{i})$ $=\delta_{\tilde{i},\tilde{k}},\ \tilde{i},\tilde{k}\in\{0,1,2\}.$ (78)
@@
In the framework of GPTs $(\mathcal{S},B)$ corresponds to a bit and $(\tilde{\mathcal{S}},\tilde{B})$ to a trit (assuming one also allows all ensembles of preparations and measurements).
@@
However in the resource theory of [44] this is not the case. We now show explicitly that there is no free operation $T$ taking $(\mathcal{S},B)$ to $(\tilde{\mathcal{S}},\tilde{B})$. Since there is only one measurement in both scenarios we drop the indices $j$ and $\tilde{j}$. Let us assume the existence of stochastic maps $q_{O}(\tilde{k}|k)$ and $q_{P}(i|\tilde{i})$ which map $B\to\tilde{B}$:
$p(\tilde{k}|\tilde{i})=\sum_{i,k=0}^{1}q_{O}(\tilde{k}|k)p(k|i)q_{P}(i|\tilde{i})$ (79)
Now using the definitions of $B$ and $\tilde{B}$ of Equation 77 we obtain:
$\delta_{\tilde{i},\tilde{k}}=\sum_{i=0}^{1}q_{O}(\tilde{k}|i)q_{P}(i|\tilde{i}).$ (80)
Let us denote by $Q_{O}$ the $3\times 2$ matrix with entries $[Q_{O}]_{\tilde{k},i}=q_{O}(\tilde{k}|i)$ and $Q_{P}$ the $2\times 3$ matrix with entries $[Q_{P}]_{i,\tilde{i}}=q_{P}(i|\tilde{i})$. Then the previous equation is equivalent to the matrix equation:
@@
This yields a contradiction since the decomposition of a $3\times 3$ matrix $A$ into matrix product $A=B\cdot C$ with $B$ a $3\times 2$ matrix and $C$ a $2\times 3$ matrix implies $C$ has rank at most 2. However $\mathbb{I}_{3}$ is a rank 3 matrix. This shows that there is no free operation $(\mathcal{S},B)\mapsto(\tilde{\mathcal{S}},\tilde{B})$ in the resource theory of [44]. Notice that the reason why in our resource theory the bit and the trit — and any other noncontextual theory — are equivalent is that our free operations include the access to the classical system. This is the origin of the difference with the resource theory of [44].
An additional contribution of the present work is the following. In [44] the authors define a notion of composite of two behaviours which is equivalent to the minimal composite of Definition 5 in our framework. They then show that this composite is consistent with their resource theory, in the sense that the free objects are closed under this composite. In the present work we show that any consistent composite of GPT systems, not just the minimal composite of Definition 5 is compatible with the resource theory of contextuality.
@@ 6 Conclusion
Generalized contextuality, a leading notion of nonclassicality, is of crucial importance both in the foundations of quantum theory and in quantum information processing. Despite this, there is no complete characterisation of generalised contextuality as a resource. In this work we address this shortcoming.
@@
A key feature of the resource theory presented is this work is that the free operations include access to classical systems. Since every noncontextual system can be univalently simulated by a classical system this entails that all noncontextual systems are equivalent in the hierarchy. Without access to free classical systems the preorder which emerges is the embeddability preorder, for which generic noncontextual systems are inequivalent. Moreover we show, by assuming that the only way of composing GPT systems with classical ones is the standard one, that any consistent composition of GPT systems is consistent with the hierarchy: if $\mathsf{A}_{1}\preceq\mathsf{B}_{1}$ and $\mathsf{A}_{2}\preceq\mathsf{B}_{2}$ then $\mathsf{A}_{1}\mathbin{\tilde{\otimes}}\mathsf{A}_{2}\preceq\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}$.
@@ Acknowledgments
The authors thank the participants of the PIMan workshop – Orange (CA), March 2019 – where the idea of this project originated. In particular, Luke Burns and Justin Dressel, who were part of the initial discussions on the project. The authors further thank Rafael Wagner for insightful explanations regarding the approach of Duarte and Amaral [44]. LC thanks Farid Shahandeh and TGa thanks Markus Müller for helpful discussions. This project started when LC was supported by the Fetzer Franklin Fund of the John E. Fetzer Memorial Trust and by the Army Research Office (ARO) (Grant No. W911NF-18-1-0178). LC also acknowledges funding from the Einstein Research Unit “Perspectives of a Quantum Digital Transformation” and from the Horizon Europe project FoQaCiA, GA no.101070558. TGa acknowledges support from the Austrian Science Fund (FWF) via project P 33730-N. TGo acknowledges support from the Austrian Science Fund. This research was funded in whole or in part by the Austrian Science Fund (FWF) via the START Prize Y1261-N. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.
@@
Appendix A Physicality of the Holevo projection
@@
In the framework of GPTs a physical map $\mathsf{A}\to\mathsf{B}$ is given by a linear map $M:V_{\mathsf{A}}\to V_{\mathsf{B}}$ such that $M(\Omega_{\mathsf{A}})\subset\Omega_{\mathsf{B}}$. $M$ can be viewed as part of a preparation procedure, preparing a state of $\mathsf{B}$ by first preparing a state $\omega_{\mathsf{A}}$ of $\mathsf{A}$ and then applying $M$ to obtain $\omega_{\mathsf{B}}=M\omega_{\mathsf{A}}$. Composing this with a measurement $\{e_{\mathsf{B}}^{i}\}_{i}$ on $\mathsf{B}$ gives outcome probabilities $e_{\mathsf{B}}^{i}(M\omega_{\mathsf{A}})$. However, one could also view this experiment as preparing the state $\omega_{\mathsf{A}}$ on system $\mathsf{A}$ and then measuring $\{e_{\mathsf{B}}^{i}\circ M\}_{i}=\{M^{*}e_{\mathsf{B}}^{i}\}_{i}$ which is a measurement on $\mathsf{A}$, hence we require that a physical map also obeys $M^{*}e_{\mathsf{B}}\in E_{\mathsf{A}}$ for all $e_{\mathsf{B}}\in E_{\mathsf{B}}$.
Definition 32 (Physical GPT map).
@@
A physical map $M$ between two system $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ and $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}},V_{\mathsf{B}})$ is given by a linear map $M:V_{\mathsf{A}}\to V_{\mathsf{B}}$ such that:
$M(\Omega_{\mathsf{A}})\subseteq\Omega_{\mathsf{B}},$ (82)
$M^{*}(E_{\mathsf{B}})\subseteq E_{\mathsf{A}}.$ (83)
It turns out that a (multivalent) simulation of $\mathsf{B}$ by $\mathsf{A}$ does not, in general, define a physical transformation $\mathsf{A}\to\mathsf{B}$. First, a simulation $(\Gamma,\Theta)$ from $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}})$ to $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}})$ is given by a pair of multivalued maps. The state simulation $\Gamma:\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}$ is not, in general, a map (since it does not assign a unique element of $\Omega_{\mathsf{B}}$ to each $\Omega_{\mathsf{A}}$). The inverse $\Gamma^{-1}$: $\Gamma(\Omega_{\mathsf{A}})\subset\Omega_{\mathsf{B}}\to\Omega_{\mathsf{A}}$ defines a map $\Gamma(\Omega_{\mathsf{A}})\to\Omega_{\mathsf{A}}$, which does not, in general, linearly extend to a map $\Gamma^{-1}:V_{\mathsf{B}}\to V_{\mathsf{A}}$ and hence cannot be a physical transformation.
Moreover, in the case where $\Gamma^{-1}$ does define a linear map $V_{\mathsf{B}}\to V_{\mathsf{A}}$ it might not meet the requirement that $(\Gamma^{-1})^{*}(E_{\mathsf{A}})\subseteq E_{\mathsf{B}}$. This shows why, within the framework of GPTs, one cannot in general associate a physical process from an underlying classical system to the GPT system for a generic ontological model.
However, there is a family of simulations which do.
@@
The HBB simulation (see Example 16) of a system $\mathsf{A}$ with finitely extremal states $\delta_{\rm ext}(\Omega_{\mathsf{A}})=X$ on the classical system $\Delta_{|X|}$ given by the multivalued maps $(\Gamma,\Theta)$ induces a physical map $\Gamma^{-1}:\Delta_{|X|}\to\mathsf{A}$.
@@
Given a GPT system $\mathsf{A}$ with finitely many extremal states $X$ the HBB model is given by the maps $(\Gamma,\Theta)$ as defined in Example 16. We first show that $\Gamma^{-1}:\Gamma(\Omega_{\mathsf{A}})\to\Omega(\mathsf{A})$ defines a map $\Delta_{|X|}\to\Omega_{\mathsf{A}}$.
The map $\Gamma^{-1}$ acts on extremal states of $\Delta_{|X|}$ as:
$\Gamma^{-1}:\delta_{x}\mapsto\omega_{x},x\in X.$ (84)
Extending the action to mixed states gives:
$\bar{\Gamma}^{-1}:\sum_{x}p_{x}\delta_{x}\mapsto\omega=\sum_{x}p_{x}\omega_{x}.$ (85)
This is consistent with $\Gamma:\omega\mapsto\{\sum_{x\in X}p_{x}\delta_{x}|p_{x}\in[0,1],\sum_{x\in X}p_{x}=1,\omega=\sum_{x\in X}p_{x}\omega_{x}\}$ since $\bar{\Gamma}^{-1}(\Gamma(\omega))=\omega$ for every $\mu\in\Gamma(\omega)$.
Since $\Delta_{|X|}$ spans $\mathbb{R}^{|X|}$ the map $\bar{\Gamma}^{-1}$ linearly extends to $\mathbb{R}^{|X|}$, defining a map $M:\mathbb{R}^{|X|}\to V_{\mathsf{A}}$.
This map satisfies $M^{*}(E_{\mathsf{A}})\subset\Delta_{m}^{*}$ which can be seen as follows:
$M^{*}(e_{\mathsf{A}})\omega_{\Delta_{m}}=e_{\mathsf{A}}(M\omega_{\Delta_{m}})=e_{\mathsf{A}}(\omega_{\mathsf{A}})\in[0,1].$ (86)
Hence $M^{*}(e_{\mathsf{A}})$ takes $\Delta_{m}$ to $[0,1]$ and, since $\Delta_{m}$ is unrestricted, $M^{*}(e_{\mathsf{A}})\in\Delta_{m}^{*}$. ∎

### v2

@@ 5.2 Relation with previous works on contextuality and GPTs
In the following section we relate the resource theory of generalized contextuality introduced in this work to previous works on generalized contextuality in the frameworks of GPTs and contrast it to the existing proposed resource theory of contextuality of [28].
@@
Other influential sources of inspiration for our work are [72, 76, 77]. As we already mentioned in the introduction, [72] introduces the notion of simplex embeddability as the geometrical criterion to assess whether a GPT system is noncontextual. [^8] As a note our notion of the canonical ontological model associated to a GPT system differs from the notion of ontological model of a GPT system of [72] since we define the preparations to be the set of ensembles of states, whereas they define the preparations to be the set of states (hence every ontological model of a GPT system is noncontextual in their work). [^9] Another work that is closely related to [72] is that of Shahandeh [78]. However, the criterion of classicality introduced there additionally requires the simplex to have the same dimension of the GPT system. We do not adopt this approach. For example, it deems the rebit stabilizer theory to be nonclassical. However, the rebit stabilizer theory, in the prepare-and-measure scenario, admits of a noncontextual ontological model given by the Spekkens toy theory [81], as has been shown in several works [83, 16, 15]. The criterion is extended to accessible GPT fragments in [76], where the latter correspond to more general mathematical objects than GPTs and characterize generic prepare-and-measure experimental setups. The work of [77] provides an algorithm for testing contextuality in any prepare-and-measure scenario. In particular, if it exists, it returns an explicit noncontextual model for the scenario and, if not, it provides the minimum amount of (depolarizing) noise which would be required until a noncontextual model would become possible. They call this measure of contextuality the robustness of nonclassicality. The latter is related to our notion of error of univalent simulation, but it is different insofar as it requires a specific noise channel (e.g. depolarizing noise channel), whereas the error of univalent simulation does not require that. Moreover, the robustness of nonclassicality is not defined as a measure within a well defined resource theory.
@@
It has been argued in [76] that in realistic experiments GPT systems are generally not appropriate to study contextuality. In particular, the following discrepancy arises: A GPT system obtained from an operational description of the experiment provides states and effects observable in practice in this experiment, but does not limit the in principle implementable ones. However, the preparation equivalences preserved by a preparation-noncontextual ontological model (see Section˜2) are in principle operational equivalences.
When studying preparation contextuality only, this discrepancy can be dealt with. Namely, one can first set-up the effect space $E_{\mathsf{A}}$ in a non-operational way, as a hypothesized set of in principle possible effects. Importantly, one has to identify each actual measurement outcome arising in the experiment with a (distribution over) the elements of $E_{\mathsf{A}}$. Their means specify a subset $\tilde{E}_{\mathsf{A}}\subseteq E_{\mathsf{A}}$ of the so-called accessible effects, which however play no role in the study of preparation contextuality. The state space $\Omega_{\mathsf{A}}$ is the set of equivalence classes of preparation procedures, but only with respect to in principle operational equivalence. That is, states that are indistinguishable by accessible effects, but could be distinguished by a hypothesized effect in $E_{\mathsf{A}}\setminus\tilde{E}_{\mathsf{A}}$ ought not be identified in $\Omega_{\mathsf{A}}$. With this set-up, one can study preparation contextuality as we discuss shortly in Section˜3.3 in terms of the GPT system $(\Omega_{\mathsf{A}},E_{\mathsf{A}})$.
@@
$\mathcal{P}=\{0,1\},\quad$ $\mathcal{M}=\{0\},\quad$ $\mathcal{D}=\{0,1\},\$$\mathcal{E}_{P}=\emptyset,\$$\mathcal{E}_{M}=\emptyset,$ (60)
$\mathcal{P}=\{0,1,2\},\quad$ $\mathcal{M}=\{0\},\quad$ $\mathcal{D}=\{0,1,2\},\quad$ $\mathcal{E}_{P}=\emptyset,\quad$ $\mathcal{E}_{M}=\emptyset.$ (61)
@@
[…]
@@
[…]
@@
[…]
Now using the definitions of $B$ and $\tilde{B}$ of Equation˜62 we obtain:
[…]
@@
[…]
An additional contribution of the present work is the following. In [28] the authors define a notion of composite of two behaviours which is equivalent to the minimal composite of Definition˜5 in our framework. They then show that this composite is consistent with their resource theory, in the sense that the free objects are closed under this composite. In the present work we show that any consistent composite of GPT systems, not just the minimal composite of Definition˜5 is compatible with the resource theory of contextuality.
@@ 6 Conclusion
Generalized contextuality, a leading notion of nonclassicality, is of crucial importance both in the foundations of quantum theory and in quantum information processing. Despite this, there is no complete characterisation of generalized contextuality as a resource. In this work we address this shortcoming.
@@
A key feature of the resource theory presented in this work is that the free operations include access to classical systems. Since every noncontextual system can be univalently simulated by a classical system this entails that all noncontextual systems are equivalent in the hierarchy. Without access to free classical systems the preorder which emerges is the embeddability preorder, for which generic noncontextual systems are inequivalent. Moreover we show, by assuming that the only way of composing GPT systems with classical ones is the standard one, that any consistent composition of GPT systems is consistent with the hierarchy: if $\mathsf{A}_{1}\preceq\mathsf{B}_{1}$ and $\mathsf{A}_{2}\preceq\mathsf{B}_{2}$ then $\mathsf{A}_{1}\mathbin{\tilde{\otimes}}\mathsf{A}_{2}\preceq\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}$.
@@ Acknowledgments
The authors thank the participants of the PIMan workshop – Orange (CA), March 2019 – where the idea of this project originated. In particular, Luke Burns and Justin Dressel, who were part of the initial discussions on the project. The authors further thank Rafael Wagner for insightful explanations regarding the approach of Duarte and Amaral [28]. LC thanks Farid Shahandeh and TGa thanks Markus Müller for helpful discussions. This project started when LC was supported by the Fetzer Franklin Fund of the John E. Fetzer Memorial Trust and by the Army Research Office (ARO) (Grant No. W911NF-18-1-0178). LC also acknowledges funding from the Einstein Research Unit “Perspectives of a Quantum Digital Transformation” and from the Horizon Europe project FoQaCiA, GA no.101070558. TGa acknowledges support from the Austrian Science Fund (FWF) via project P 33730-N. This research was partly funded by the Austrian Science Fund (FWF) 10.55776/PAT2839723. TGo acknowledges support from the Austrian Science Fund. This research was funded in whole or in part by the Austrian Science Fund (FWF) via the START Prize Y1261-N. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.
@@
Appendix A On the physicality of simulations
@@
In the framework of GPTs a physical map $\mathsf{B}\to\mathsf{A}$ is given by a linear map $M:V_{\mathsf{B}}\to V_{\mathsf{A}}$ satisfying $(\Omega_{\mathsf{B}})\subseteq\Omega_{\mathsf{A}}. MM$ can be viewed as part of a preparation procedure, preparing a state of $\mathsf{A}$ by first preparing a state $\omega_{\mathsf{B}}$ of $\mathsf{B}$ and then applying $M$ to obtain $\omega_{\mathsf{A}}$ given by $M(\omega_{\mathsf{B}})$. Composing this with a measurement $\{e_{\mathsf{A}}^{i}\}_{i}$ on $\mathsf{A}$ gives outcome probabilities $e_{\mathsf{A}}^{i}\cdot M(\omega_{\mathsf{B}})$. However, one could also view this experiment as preparing the state $\omega_{\mathsf{B}}$ on system $\mathsf{B}$ and then measuring $\{e_{\mathsf{A}}^{i}\circ M\}_{i}=\{M^{*}e_{\mathsf{A}}^{i}\}_{i}$ which is a measurement on $\mathsf{B}$, hence we require that a physical map also obeys $M^{*}(e_{\mathsf{A}})\in E_{\mathsf{B}}$ for all $e_{\mathsf{A}}\in E_{\mathsf{A}}$.
Definition 23 (Physical GPT transformation).
@@
A physical GPT transformation $M$ between two GPT systems, from $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}},V_{\mathsf{B}})$} to $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$, is given by a linear map $M:V_{\mathsf{B}}\to V_{\mathsf{A}}$ satisfying
$M(\Omega_{\mathsf{B}})\subseteq\Omega_{\mathsf{A}}\qquad\text{and}\qquad M^{*}(E_{\mathsf{A}})\subseteq E_{\mathsf{B}}.$ (67)
One may wonder whether an exact simulation $(\Gamma,\Theta)$ of a system $\mathsf{A}$ by a system $\mathsf{B}$ can correspond to a physical transformation $M:\mathsf{B}\to\mathsf{A}$, [^11] There is another possibility: namely, the simulation may correspond to a physical transformation $W:\mathsf{A}\to\mathsf{B}$, in the sense that the state simulation $\Gamma$ assigns, to each state $\omega\in\Omega_{\mathsf{A}}$, the set of those states of $\mathsf{B}$ that could give rise to $\omega$ when $M$ is applied. We do not treat this case here. in the sense that the state simulation $\Gamma$ assigns, to each state $\omega\in\Omega_{\mathsf{A}}$, a subset of those states of $\mathsf{B}$ that could give rise to $\omega$ when $M$ is applied.
Definition 24 (Physical realisation of a simulation).
Consider a simulation $(\Gamma,\Theta)$ of a GPT system $\mathsf{A}$ by a GPT system $\mathsf{B}$ and a physical GPT transformation $M:\mathsf{B}\to\mathsf{A}$. We say that $M$ is a physical realisation of $(\Gamma,\Theta)$ iff we have $\Gamma(\omega)\subseteq M^{-1}(\omega)$ for all $\omega\in\Omega_{\mathsf{A}}$ and $\Theta(e)\supseteq\{M^{*}(e)\}$ for all $e\in E_{\mathsf{A}}$. [^12] Here, $M^{-1}(\omega)$ refers to the preimage of $\omega$ under $M$ and $\{M^{*}(e)\}$ refers to the singleton set whose element is $M^{*}(e)$.
[…]
Requiring a simulation $(\Gamma,\Theta)$ of a GPT system $\mathsf{A}$ by a system $\mathsf{B}$ to admit of a physical realisation is a very strong condition. It has the following consequences.
Proposition 25.
Consider a simulation $(\Gamma,\Theta)$ of a system $\mathsf{A}$ by a system $\mathsf{B}$ with a physical realisation $M:\mathsf{B}\to\mathsf{A}$.
  1. 1.
    $M$ is surjective on states, i.e. we have $M(\Omega_{\mathsf{B}})=\Omega_{\mathsf{A}}$.
  2. 2.
    If $\dim(V_{A})=\dim(V_{B})$ holds, then $(\Gamma,\Theta)$ is a univalent simulation and $M$ is an isomorphism.
Proof.
By Definition˜24, we have $\Gamma(\omega)\subseteq M^{-1}(\omega)$ for all $\omega\in\Omega_{\mathsf{A}}$. In particular, the preimage of every such $\omega$ under $M$ must be non-empty, which means $M(\Omega_{\mathsf{B}})\supseteq\Omega_{\mathsf{A}}$. Combining this with the first condition in (67) gives $M(\Omega_{\mathsf{B}})=\Omega_{\mathsf{A}}$.
Let us now assume that the underlying vector spaces of $\mathsf{A}$ and $\mathsf{B}$ have the same dimension given by $d\in\mathbb{N}$. We prove that $\Gamma$ is single-valued by contradiction. To this end, assume that there are two distinct $\gamma_{1},\gamma_{2}\in\Gamma(\omega)$ for some $\omega\in\Omega_{\mathsf{A}}$. By empirical adequacy of the simulation, we must have
$f\cdot\gamma_{1}=f\cdot\gamma_{2}$ (69)
for all $f\in\Theta(E_{\mathsf{A}})$. In particular, this means that the span of $\Theta(E_{\mathsf{A}})$ has dimension strictly less than $d$ and no effect in $\Theta(E_{\mathsf{A}})$ can distinguish two states of $\Omega_{\mathsf{B}}$ that differ by a scalar multiple of $\gamma_{1}-\gamma_{2}$.
By the fact that effects of $\mathsf{A}$ can be distinguished by its states, and the fact that $\Omega_{\mathsf{A}}$ is a convex set (not containing the zero vector), there are $d$ states $\rho_{i}\in\Omega_{\mathsf{A}}$ whose span is the whole of $V_{\mathsf{A}}$. To distinguish these $d$ states, one needs at least $d$ linearly independent effects, let us denote a choice of these by $e_{i}\in E_{\mathsf{A}}$. By empirical adequacy, any choice of $d$ states $\sigma_{i}\in\Gamma(\rho_{i})$ of $\mathsf{B}$ must also span $V_{\mathsf{B}}$ and be distinguishable by effects in
$\Theta\left(\{e_{i}\}_{i=1}^{d}\right)\subseteq\Theta(E_{\mathsf{A}}).$ (70)
This is now a contradiction because $\Theta(E_{\mathsf{A}})$, as we established earlier, has dimension strictly less than $d$. In conclusion, $\Gamma$ must be single-valued. An analogous argument can be used to show that $\Theta$ must also be single-valued.
By the univalence of the simulation, we can thus think of $\Gamma$ and $\Theta$ as convex-linear maps $\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}$ and $E_{\mathsf{A}}\to E_{\mathsf{B}}$ respectively. Since $\mathrm{span}(\Omega_{\mathsf{A}})=V_{\mathsf{A}}$ and $0\not\in\Omega_{\mathsf{A}}$ (by the second condition in (5)) it follows that we can actually identify the state simulation with a linear map $V_{\mathsf{A}}\to V_{\mathsf{B}}$ denoted by $\bar{\Gamma}$.
Since $\Gamma$ is injective by Implication (20) (which implies that $\bar{\Gamma}$ is as well) and we have $V_{\mathsf{A}}\cong V_{\mathsf{B}}$ by assumption, $\bar{\Gamma}$ is an isomorphism.
By Definition˜24, we then have $M\Gamma(\omega)=\omega$ for all $\omega\in\Omega_{\mathsf{A}}$. By linearity, we can extend this equation to the span of $\Omega_{\mathsf{A}}$ to get $M\bar{\Gamma}=\mathrm{id}_{V_{\mathsf{A}}}$, which means that $M$ is the inverse of $\bar{\Gamma}$ since the latter is an isomorphism. Thus, $M$ itself must be an isomorphism. ∎
In the case of unequal dimensions of $V_{\mathsf{A}}$ and $V_{\mathsf{B}}$, a simulation can indeed be realised also by a physical GPT transformation that is not an isomorphism. One example is the simulation of the n-level classical system $\Delta_{n}$ by the quantum one $\mathsf{Q}_{n}$ mentioned above. The completely dephasing map from (68) gives a realisation thereof.
One of the consequences of Proposition˜25 is that whenever there exists a simulation of a GPT system $\mathsf{A}$ by $\mathsf{B}$ and $\dim(V_{\mathsf{A}})=\dim(V_{\mathsf{B}})$ holds, the state spaces of $\mathsf{A}$ and $\mathsf{B}$ must be isomorphic as convex sets. However, the same is not true for the effect spaces, as the following example shows.
Example 26 (Physical realisation of univalent simulation for inequivalent effect spaces).
Consider two systems $\mathsf{A}$ and $\bar{\mathsf{A}}$ with identical underlying vector spaces, where we also have $\Omega_{\mathsf{A}}=\Omega_{\bar{A}}$ and $E_{\bar{A}}\subsetneq E_{\mathsf{A}}$. Then the inclusion maps on states and effects provide a univalent simulation of $\bar{\mathsf{A}}$ by $\mathsf{A}$. Moreover, it is realisable by the physical GPT transformation given by the identity $\mathbb{I}:V_{\mathsf{A}}\to V_{\bar{\mathsf{A}}}$}, since we have $\mathbb{I}(\Omega_{\mathsf{A}})=\Omega_{\bar{\mathsf{A}}}$ and $\mathbb{I}^{*}(E_{\bar{\mathsf{A}}})\subseteq E_{\mathsf{A}}$. Even though the identity is invertible as a linear map, its inverse $\mathbb{I}:V_{\bar{\mathsf{A}}}\to V_{\mathsf{A}}$ is not a physical GPT transformation since $\mathbb{I}^{*}(E_{\mathsf{A}})\subseteq E_{\bar{\mathsf{A}}}$ is false by assumption.
We can apply Proposition˜25 to the (convexified) Spekkens toy bit $\mathsf{T}_{2}$ (with six extremal states), which has a non-contextual ontological model, i.e. it can be simulated by a classical GPT system, in particular by $\Delta_{4}$ with four extremal states. Since these two GPT systems have isomorphic underlying vector spaces and non-isomorphic state spaces, it follows that there can be no physical realisation $\Delta_{4}\to\mathsf{T}_{2}$ of the univalent simulation of $\mathsf{T}_{2}$ by $\Delta_{4}$.
Interestingly, for any GPT system $\mathsf{A}$ with finitely many extremal states there is an ontological model thereof, which admits of a physical realisation.
@@
The HBB simulation (see Example˜9) of a GPT system $\mathsf{A}$ with $n\in\mathbb{N}$ extremal states by the classical system $\Delta_{n}$ has a physical realisation.
@@
Given a GPT system $\mathsf{A}$ with a finite set $\{\omega_{i}\}_{i=1}^{n}$ of extremal states, the HBB model is given by the maps $(\Gamma,\Theta)$ as defined in Example˜9. We first show that the opposite $\Gamma^{-1}:\Gamma(\Omega_{\mathsf{A}})\to\Omega_{\mathsf{A}}$} of $\Gamma$ (which is a function as shown in (20)) uniquely extends to a linear map $M:\mathbb{R}^{n}\to V_{\mathsf{A}}$.
The map $\Gamma:\Omega_{\mathsf{A}}\to\Delta_{n}$ is a multivalued function defined as:
$\Gamma(\omega)=\left\{\sum_{i=1}^{n}p_{i}\delta_{i}\nonscript\;\,\middle|\allowbreak\nonscript\;\,\mathopen{}p_{i}\in[0,1],\,\sum_{i=1}^{n}p_{i}=1,\,\omega=\sum_{i=1}^{n}p_{i}\omega_{i}\right\}.$ (71)
For an arbitrary element $\delta\in\Delta_{n}$, which can be uniquely decomposed as
$\delta=\sum_{i}p_{i}\delta_{i},$ (72)
the function $\Gamma^{-1}$ is given by
$\Gamma^{-1}\left(\delta\right)=\sum_{i}p_{i}\omega_{i}$ (73)
and is in particular a convex-linear map $\Delta_{n}\to\Omega_{\mathsf{A}}$.
Since $\Delta_{n}$ spans $\mathbb{R}^{n}$ and does not contain the origin, the map $\Gamma^{-1}$ uniquely extends to a linear map $M:\mathbb{R}^{n}\to V_{\mathsf{A}}$ satisfying $M(\Delta_{n})=\Omega_{\mathsf{A}}$ and $\Gamma(\omega)\subseteq M^{-1}(\omega)$ for all $\omega\in\Omega_{\mathsf{A}}$.
This map also satisfies $M^{*}(E_{\mathsf{A}})\subseteq\Delta_{n}^{*}$, since for any $e\in E_{\mathsf{A}}$ we have
$\forall\delta\in\Delta_{n}\quad\quad M^{*}(e)\cdot\delta=e\cdot M(\delta)\in[0,1],$ (74)
so that $M^{*}(e)$ is an element of $\Delta_{n}^{*}$. Therefore, $M$ is a physical GPT transformation.
The univalent effect simulation map is given (as a function) by $\Theta(e)=\sum_{i}(e\cdot\omega_{i})\delta_{i}^{*}$. We now have, for any $j\in\{1,2,\ldots,n\}$},
$M^{*}(e)\cdot\delta_{j}=e\cdot M(\delta_{j})=e\cdot\omega_{j}=\sum_{i}(e\cdot\omega_{i})(\delta_{i}^{*}\cdot\delta_{j})=\Theta(e)\cdot\delta_{j},$ (75)
so that $M^{*}(e_{\mathsf{B}})=\Theta(e_{\mathsf{B}})$ holds by (8). In conclusion, this shows that $M$ is indeed a physical realisation of the HBB simulation. ∎